Mission Operations Days: 2004/318 to 2004/324
November 13, 2004 00:00:00 GMT to November 19, 2004 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|AM-1 "Terra" is in Normal Mode|
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
|Blackbody||A On (290K); B Off||Nominal Science Mode|
|Calibration Electronics||A On; B Off||Nominal Science Mode|
|Control Processor||A On; B Off||Nominal Science Mode|
|Nadir Door||Unlatched, Open||Nominal Science Mode|
|Space View Door||Unlatched, Open||Nominal Science Mode|
|Solar Diffuser Door||Unlatched, Screened||Permanent Configuration|
|FDDI Formatter||A On; B Off||Nominal Science Mode|
|FDDI Port||A On; B Off||Nominal Science Mode|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal Science Mode|
|Format Processor||A Off; B On||Nominal Science Mode|
|Power Supplies||#1 On, #2 Off||Nominal Science Mode|
|PV VIS FPA||A On; B Off||Nominal Science Mode|
|PV NIR FPA||A On; B Off||Nominal Science Mode|
|PV SMIR FPA||A On; B Off||Nominal Science Mode|
|PV LWIR FPA||A On; B Off||Nominal Science Mode|
|PC LWIR FPA||A On; B Off||Nominal Science Mode|
|RC Outgas Heaters||All 3 (CS, IS, OS) Off||Nominal Science Mode|
|RC LWIR FPA Heater||On (83K)||Nominal Science Mode|
|RC SMIR FPA Heater||Off||Nominal Science Mode|
|Scan Assembly||A On; B Off||Nominal Science Mode|
|SDSM||Off||Nominal Science Mode|
|SRCA||Off||Nominal Science Mode|
|PS1 Survival Heater||Enabled||Nominal Science Mode|
|PS2 Survival Heater||Enabled||Nominal Science Mode|
|Timing Generator||A On; B Off||Nominal Science Mode|
|Flight Software||Rev BD + 11 sets of patches||Nominal Science Mode|
|Inhibit Ids Set||none||Nominal Science Mode|
|TMON 66||Disabled||Nominal Science Mode|
|TMON 67||Disabled||Nominal Science Mode|
|TMON 68||Enabled||Nominal Science Mode|
|TMON 69||Disabled||Nominal Science Mode|
This Week’s Completed MODIS Activities:
Saturday, November 13, 2004
Sunday, November 14, 2004
Monday, November 15, 2004
Tuesday, November 16, 2004
2004/321 20:29:43-20:38:31 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/321 21:10:00-21:10:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/321 21:19:02-21:54:12 ATC MOD_OA19_SRCA_FULL_RAD_A.6 OA-19: SRCA Full Radiometric Calibration.
Data collection from 21:34:56-21:53:13
Wednesday, November 17, 2004
Thursday, November 18, 2004
Friday, November 19, 2004
This Week’s Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Calibrations for November will include biweekly SDSM, monthly SRCA Full and 10W Radiometric, and bimonthly SRCA Full Spatial calibrations.
Saturday, November 20, 2004
Extra 20 minutes of DAY rate during ground track night
2004/325 20:13:09-20:50:37 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate
2004/325 20:50:37-21:10:00 ATC MOD_OA04_NIGHT_RATE.1 OA-04: Formatter to Night Rate (10 min before groundtrack night)
2004/325 21:10:00-21:30:00 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate (night ground track)
2004/325 21:30:00-22:00:03 ATC MOD_OA04_NIGHT_RATE.1 OA-04: Formatter to Night Rate (return to normal night rate)
2004/325 22:00:03-22:41:32 ATC MOD_OA03_DAY_RATE.2 OA-03: Formatter to Day Rate (10 min after groundtrack day)
Monday, November 22, 2004
2004/327 17:30:00-17:30:31 ATC MOD_SRCA_14_6_4_1_SETUP.1 Turns on the SRCA and sets it's lamp config. to use 10W lamp #2
2004/327 17:38:51-19:30:16 ATC MOD_SRCA_SPECIAL_10W_OPEN.2 Performs SRCA 10W continuous radiometric activity with the filter in the OPEN position
Data collect from 17:39:55-19:30:16
Tuesday, November 23, 2004
2004/328 08:00:00-08:00:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/328 08:07:26-09:05:37 ATC MOD_OA23_SRCA_FULL_SPAT_A.4 OA-23: SRCA Full Spatial Calibration - A Side
Data collection from 08:23:20-09:04:32.
S/C Night from 08:30:20-09:04:13.
The December Lunar Calibration Roll will be on day 2004/336 (December 1, 2004). Data collect will be from 21:17:44-21:24:30 GMT with a roll angle of –8.6983 degrees.
MODIS NAD (Nadir Aperture Door) Anomaly:
Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.
General Instrument Comments:
MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.
MODIS Telemetry Trends:
Nominal telemetry trending is ongoing.
Non-MODIS Significant Events:
Terra Solid State Recorder Anomaly
The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA’s that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous report am1weekly_2003-305_2003-311.doc. The Terra FOT is currently evaluating plans to perform a power recycling of the SSR in an attempt to regain functionality of the “bad” PWAs. This activity will be performed at a time when it is least detrimental to critical science data.
The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.
Limited Life Item Status:
SRCA 10W Lamp #1: 237.6 of 500 hours
SRCA 10W Lamp #2: 172.4 of 500 hours
SRCA 10W Lamp #3: 182.2 of 500 hours
SRCA 10W Lamp #4: 66.8 of 500 hours
SRCA 1W Lamp #1: 567.0 of 4000 hours
SRCA 1W Lamp #2: 278.7 of 4000 hours
Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements